in a cricket match bye runs were double of the no-balls,the remaining score was greater by six than twelve times the number of bye-runs.if the total score was 276.how were the runs obtained?

Answer by mananth(16946)

No balls----------x
byes = 2x
total runs = 276
remaining runs = (276-3x)=12*2x+6

276-3x=24x+6
27x=270
x=10
No ball
s = 10
Byes = 20
Runs = 246